Please answer -3,-1,1,3,5 using recursive formula
Answer by Theo(13342)

if you start with -3 and keep adding 2 to it, you will wind up with 5.

-3 + = -1 + 2 = 1 + 2 = 3 + 2 = 5

your recursive formula would be A[n+1] = A[n] + 2

you would need a starting point, such as A[1] = -3

you would run your recursive program 4 times.

A[2] = A[1] + 2 = -3 + 2 = -1
A[3] = A[2] + 2 = -1 + 2 = 1
A[4] = A[3] + 2 = 1 + 2 = 3
A[5] = A[4] + 2 = 3 + 2 = 5

your result would be A[5] = 5